# Long-horizon Locomotion and Manipulation on a Quadrupedal Robot with Large Language Models

### v1

@@ IV-B Ablation studies
success rate $\uparrow$ normalized distance $\downarrow$
@@
(a) Light-switching task.
(b) Package-delivery task.
Figure 5: Execution of LLM-generated long-term strategies for two benchmark tasks in the simulation and real world.
(a) Quadrupedal locomotion controller.
(b) Bipedal manipulation controller.
(c) Object-pushing motion planner.
(d) Obstacle-avoidance motion planner.
Figure 6: Visualization of versatile behaviors driven by the RL-based skill repertoire, covering short-term locomotion and manipulation controllers (a)(b) and mid-term motion planning skills (c)(d). Each policy is capable of a range of continuously parameterized problems.
IV-C Main results
We test our system on two aforementioned benchmark tasks in the real world. As shown in Fig. 4(a), we execute the generated code together with the RL motion skills to turn off the light. In the initial setup, the height of the light button is out of reach with any single skill, and the first level of stairs is too high for the robot to climb directly. Considering the scene configuration, the LLM planner decides to push a lower box in front of the existing stairs so that the robot can use it as an intermediate platform to climb up then to touch the button. Finally, it runs a sit-down policy to transit to the quadrupedal pose. In Fig. 4(b), we show how the quadrupedal robot delivers a package into a room with the door closed initially. To get access to the room, our reasoning layer figures out a strategy to ring a bell to notify the human inside the room to open the door. To protect the package from unnecessary collisions, it leverages the obstacle-avoidance locomotion skill when moving to the bell and before pushing the package.
IV-D More analysis
Analysis of low-level RL skills: We showcase how the parameterized motion planning and controlling policies enable versatile motions in Fig. 6. The quadrupedal locomotion policy can walk or climb up boxes of various sizes. Fig. 5(a) shows snapshots when climbing the height of 0.35m and 0.15m. The bipedal locomotion and manipulation control policy could touch target spots with different relative positions to the robot, as shown in Fig. 5(b), where the robot is adapting itself when the target position is lower, higher, to its left or right. As for the motion planning policies, the object-pushing policy demonstrates some adjustment trajectories when aligning the pose of the box to the target, shown in Fig. 5(c). We test the robot to walk to different goal positions while avoiding collision with an obstacle shown in Fig. 5(d), and find the policy can choose different paths according to the target position.
Figure 7: One typical failure case in the light-switching task. The quadruped robot falls to the ground from the stairs after it loses balance during the execution of the bipedal manipulation skill.
Failure case: The quadrupedal robot may enter dead ends when the execution of some low-level skills fails. As shown in Fig. 7, the robot occasionally loses balance after standing up and then falls from the stairs. Since our system gives plans in an open-loop manner currently, the robot cannot get an updated plan in such cases. In addition, our system cannot invent policies beyond the fixed set of low-level skills, and thus cannot recover from fatal states without a reset policy.
@@ V Conclusion
We present an LLM-based system for tackling long-horizon tasks with a quadruped robot that extends the autonomy of quadrupeds from shorter-term motions for specific skills to long-range complex behaviors that require orchestrating multiple locomotion and manipulation policies. The high-level task planning is enabled with our design of multiple LLM agents for decomposing the whole task into robot actions and then grounded to parameterized robot API calls. The low-level control that exploits the agility of quadrupeds to unlock rich environment interactions is instantiated with model-free RL. Our LLM reasoning layer plans in an open loop. Allowing it to improve from feedback such as execution traces with in-context learning is an interesting future direction. Currently, the reasoning layer plans over a fixed set of low-level robot skills crafted by experts. Future research could investigate the possibility of actively discovering new skills to learn when encountering unprecedented circumstances.
@@ Acknowledgements
Z. L. and K. S. acknowledge financical support from InnoHK of the Government of the Hong Kong Special Administrative Region via the Hong Kong Centre for Logistics Robotics. The authors thank Xiaomi Inc. for providing CyberDog2 for experiments.

### v2

@@ IV-B Ablation study analysis
Success rate $\uparrow$ NormDist $\downarrow$
@@
Ablation on fine-tuning RL skills for robust chaining: We compare with a variant without fine-tuning the low-level skills from augmented initial states in the light switching task. As reported in Table II, our method with additional fine-tuning performs significantly better. Since the two methods use the same high-level plans, it indicates that the proposed fine-tuning stage is critical to the robustness of the system when transiting between different skills.
IV-C Real world experiments
To facilitate the successful deployment of RL policies to a real Xiaomi Cyberdog2 [39] robot, we add domain randomization in motion control and observation noise during training in simulation. We attach multiple AprilTags [40] to objects and use a forward-view RealSense D430 camera on the robot for pose estimation. We also randomly freeze the observations to mimic the situation when the detection fails due to the motion blur of captured images and the loss of target visibility.
We test our system on two aforementioned benchmark tasks in the real world. As shown in the upper part of Fig. 5, we execute the generated code together with the RL motion skills to turn off the light. In the initial setup, the height of the light button is out of reach with any single skill, and the first level of stairs is too high for the robot to climb directly. Considering the scene configuration, the LLM planner decides to push a lower box in front of the existing stairs so that the robot can use it as an intermediate platform to climb up then to touch the button. Finally, it runs a sit-down policy to transit to the quadrupedal pose. We also show how the quadrupedal robot delivers a package into a room with the door closed initially. To get access to the room, our reasoning layer figures out a strategy to ring a bell to notify the human inside the room to open the door. To protect the package from unnecessary collisions, it leverages the obstacle-avoidance locomotion skill when moving to the bell and before pushing the package.
Failure case: The robot may enter dead ends when the execution of some low-level skills fails. Since our system gives plans in an open-loop manner, the robot cannot get an updated plan in such cases. In addition, our system cannot invent policies beyond the fixed set of low-level skills, and thus cannot recover from fatal states without a reset policy.
@@ V Conclusion
In summary, we present an LLM-based system for tackling long-horizon tasks with a quadruped robot that extends the autonomy of quadrupeds from shorter-term motions for specific skills to long-range complex behaviors that require orchestrating multiple locomotion and manipulation policies. The high-level task planning is enabled with our design of multiple LLM agents for decomposing the whole task into robot actions and then grounded to parameterized robot API calls. The low-level control that exploits the agility of quadrupeds to unlock rich environment interactions is instantiated with model-free RL. Our LLM reasoning layer plans in an open loop. Allowing it to improve from feedback such as execution traces with in-context learning is an interesting future direction. Currently, the reasoning layer plans over a fixed set of low-level robot skills crafted by experts. Based on this work, exciting future research could investigate the possibility of actively discovering new skills to learn when faced with unprecedented circumstances.
@@ Acknowledgements
Yi Wu is supported by 2030 Innovation Megaprojects of China(Programme on New Generation Artificial Intelligence) Grant No. 2021AAA0150000. Z. L. and K. S. acknowledge financical support from InnoHK of the Government of the Hong Kong Special Administrative Region via the Hong Kong Centre for Logistics Robotics. The authors thank Xiaomi Inc. for providing CyberDog2 for experiments. The authors are grateful to Mengdi Xu for the insightful discussions that significantly contributed to this work.